OpenStudy (anonymous):

which is a table of values for y=x-6

OpenStudy (anonymous):

A

OpenStudy (anonymous):

B

OpenStudy (anonymous):

C

OpenStudy (anonymous):

D

OpenStudy (gabylovesyou):

fill in the values.. Lets try the first choice... you ave y = x - 6 1 = -5 - 6 Is this true ?

OpenStudy (gabylovesyou):

what is -5 - 6 ?

OpenStudy (anonymous):

-11 I THINK

OpenStudy (gabylovesyou):

exactly... its not 1 so choice A is NOT CORRECT.

OpenStudy (anonymous):

ok

OpenStudy (gabylovesyou):

Lets try next one... can you substitute the values for me ? y = x - 6

OpenStudy (anonymous):

oh so its earthier b or c

OpenStudy (gabylovesyou):

exactly. fill in the values for me..

OpenStudy (gabylovesyou):

y = x - 6 what is y and what is x ?

OpenStudy (anonymous):

so y is -11

OpenStudy (gabylovesyou):

-11 = x - 6 what is x ?

OpenStudy (anonymous):

x is -5

OpenStudy (gabylovesyou):

-11 = -5 - 6 this is correct.. BUT lets keep going to the others and substitute.

OpenStudy (anonymous):

ok

OpenStudy (gabylovesyou):

-2 = -8 - 6 ?? Is this correct ?

OpenStudy (gabylovesyou):

Does -8 - 6 EQUAL to -2 ?

OpenStudy (anonymous):

no its -14 i think

OpenStudy (gabylovesyou):

exactly.. so choice B is incorrect. That leaves you with ???? Whats ur answer ?

OpenStudy (anonymous):

c

OpenStudy (gabylovesyou):

good job :)

OpenStudy (anonymous):

thx